3. Stock A earned -2% last year and 34% this year (each return is equally likely). Stock B earned 6% last year and 2% this year (each return is equally likely). For Stock B: average is 4%, variance is .0004 and standard deviation is 2%. If you put 60% Stock A and 40% in Stock B, find the return and variance of the portfolio.
Answer #1

Return of the portfolio
Last year=60%*(-2%)+40%*6%=1.200%
This year=60%*34%+40%*2%=21.200%

Return=(1.2%+21.2%)/2=11.20%

Variance=((1.2%-11.2%)^2+(21.2%-11.2%)^2)/(2)=0.01